# SWI: Speaking with Intent in Large Language Models

### v1

@@ Abstract
Intent, typically clearly formulated and planned, functions as a cognitive framework for reasoning and problem-solving. This paper introduces the concept of Speaking with Intent (SWI) in large language models (LLMs), where the explicitly generated intent encapsulates the model’s underlying intention and provides high-level planning to guide subsequent analysis and communication. By emulating deliberate and purposeful thoughts in the human mind, SWI is hypothesized to enhance the reasoning capabilities and generation quality of LLMs. Extensive experiments on mathematical reasoning benchmarks consistently demonstrate the superiority of Speaking with Intent over Baseline (i.e., generation without explicit intent). Moreover, SWI outperforms answer-trigger prompting methods Chain-of-Thought and Plan-and-Solve and maintains competitive performance with the strong method ARR (Analyzing, Retrieving, and Reasoning). Additionally, the effectiveness and generalizability of SWI are solidified on reasoning-intensive question answering (QA) and text summarization benchmarks, where SWI brings consistent improvement to the Baseline generation. In text summarization, SWI-generated summaries exhibit greater accuracy, conciseness, and factual correctness, with fewer hallucinations. Furthermore, human evaluations verify the coherence, effectiveness, and interpretability of the intent produced by SWI. This proof-of-concept study creates a novel avenue for enhancing LLMs’ reasoning abilities with cognitive notions. [^1] Appendix A Reproducibility Statement. Source code: https://github.com/YuweiYin/SWI
@@ 1 Introduction
Large language models (LLMs) (Zhao et al., 2023; Min et al., 2023; Minaee et al., 2024) have revolutionized Natural Language Processing (NLP) with their excellent generative capabilities (Hurst et al., 2024; Anthropic, 2024; Team et al., 2024). Many recent benchmarks focus on evaluating complex reasoning in mathematics problems (Cobbe et al., 2021; Hendrycks et al., 2021b; Lightman et al., 2024; Vendrow et al., 2025) and multidisciplinary questions (Hendrycks et al., 2021a; Srivastava et al., 2023; Suzgun et al., 2023; Wang et al., 2024e), presenting significant challenges for LLMs. Enhancing LLM reasoning is vital for their ongoing development (Huang & Chang, 2022; Qiao et al., 2023; Patil, 2025).
Recent advancements have introduced various approaches to improve LLM reasoning (Yu et al., 2024; Cheng et al., 2025; Liu et al., 2025a; Xu et al., 2025). Chain-of-Thought (CoT) prompting enhances reasoning (Li et al., 2024b) by providing step-by-step solutions in few-shot exemplars (Wei et al., 2022) or explicitly instructing the model to “think step by step” (Kojima et al., 2022). Beyond CoT variants (Yao et al., 2023a; Wang et al., 2023; Yasunaga et al., 2024), Yin & Carenini (2025) propose Analyzing, Retrieving, and Reasoning (ARR), an answer-trigger prompting method that systematically guides LLMs in question answering and outperforms CoT prompting.
@@
Inspired by ARR (Yin & Carenini, 2025), where intent analysis is identified as its most effective component, we further investigate the powerful potential that intent can unleash in enhancing LLM reasoning. Intent, the goal-oriented intention in our mind (Adams, 1986; Mele, 1989; Mele & Moser, 1994), serves as a guiding framework for problem-solving. As illustrated in Figure 1(a), human reasoning (System 2 (Kahneman, 2011)) typically follows a structured thinking loop where intent directs problem analysis and logical reasoning. Hence, we hypothesize that enabling LLMs to explicitly speak with their own intent—rather than merely analyzing the intent of questions—can replicate this meta-cognitive planning process, thereby improving the reasoning ability and generation quality of LLM.
@@
Figure 1: SWI Overview. (a) The intent, functioning as meta-thought and planning, guides the analysis with reasoning to answer the question. (b) A concrete example generated by our method: Speaking with Intent (SWI) in large language models (LLMs). (c) The performance improvement on mathematical reasoning and multiple-choice QA tasks brought by SWI.
@@
This work introduces Speaking with Intent (SWI), requiring LLMs to articulate their own intent as a planning mechanism during generation. Due to the autoregressive nature of LLMs (Radford et al., 2019) and the attention mechanism (Vaswani et al., 2017), explicitly stated intent provides high-level guidance for subsequent analysis and reasoning. Specifically, we implement SWI using instruction-following LLMs (Ouyang et al., 2022; Rafailov et al., 2023), incorporating system and user prompts for the model (Dubey et al., 2024) to generate intent. As shown in Figure 1(b), each analytical step (in blue) in solving a math problem is guided by a preceding intent statement (in orange), which is a freely generated text instead of a predefined class as in traditional intent modeling (Weld et al., 2022).
@@
To comprehensively evaluate the effectiveness and generalizability of the proposed SWI method, we conduct experiments across three diverse task categories: mathematical reasoning, multiple-choice QA, and text summarization. As shown in Figure 1(c), LLM speaking with intent (SWI) consistently outperforms Baseline, which generates responses without explicit intent. Furthermore, SWI surpasses answer-trigger prompting methods such as Chain-of-Thought (CoT) (Kojima et al., 2022) and Plan-and-Solve (PS) (Wang et al., 2023). Compared with the strong method ARR (Analyzing, Retrieving, and Reasoning) (Yin & Carenini, 2025), SWI maintains competitive performance on mathematical reasoning tasks. For text summarization, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations (Ji et al., 2023; Li et al., 2024a) in the output.
Additionally, we perform human evaluations to assess the coherence, effectiveness, and interpretability of the intent generated by our SWI method. Evaluators largely agree on the quality of the generated intent, particularly for mathematical reasoning tasks. The evaluation results confirm that Speaking with Intent in LLMs enhances task performance and output explainability. The key contributions of this work are as follows:
  • 1.
    This paper introduces Speaking with Intent (SWI) in LLMs, where the generated intent effectively guides problem analysis and logical reasoning, improving performance across various benchmarks. The source code will be publicly available.
  • 2.
    Extensive experiments across diverse reasoning and generation tasks, including mathematical reasoning, multiple-choice QA, and text summarization, demonstrate that SWI consistently outperforms the Baseline and surpasses established answer-trigger prompting methods such as Chain-of-Thought.
  • 3.
    Human evaluations validate the coherence, effectiveness, and interpretability of the intent generated by SWI. Rather than intent classification, our evaluation practice provides standards for assessing freely generated intent statements.
2 Related Work
LLM Reasoning.
Recent LLM research increasingly focuses on enhancing reasoning abilities (Sun et al., 2023; Patil, 2025; Cheng et al., 2025; Liu et al., 2025a). Chain-of-Thought (CoT) prompting facilitates problem-solving by prompting LLMs to generate intermediate reasoning steps (Kojima et al., 2022; Wei et al., 2022; Li et al., 2024b; Yeo et al., 2025). Building on CoT, various reasoning techniques have emerged (Xu et al., 2025). Among them, a recent research ARR (Yin & Carenini, 2025) consistently outperforms CoT on multiple QA tasks, where intent analysis of questions is its most effective component. Inspired by ARR, we enable LLMs to articulate their intent, using it to guide subsequent analysis and reasoning for improved task performance. Results show that Speaking with Intent (SWI) surpasses CoT and ARR on multiple benchmarks. As an intent-driven approach, SWI introduces a novel framework for advancing LLM reasoning and cognitive capabilities.
LLM Planning.
Planning plays a critical role in advancing artificial intelligence (Wang et al., 2024c), particularly in developing LLM agents in real and complex environments (Qiao et al., 2024; Mialon et al., 2024; Xu et al., 2024; Drouin et al., 2024; Boisvert et al., 2024; Xie et al., 2024a; b). Aside from improving the tool-use capabilities (Wang et al., 2024f; Wu et al., 2024; Huang et al., 2024; Wang et al., 2024a; Liu et al., 2025b), recent LLM planning methods enhance reasoning abilities (Huang et al., 2022; Yao et al., 2023b; Hao et al., 2023; Valmeekam et al., 2023; Jiao et al., 2024) by generating problem-solving plans explicitly (Wang et al., 2023; Li & Zhang, 2024; Lyu et al., 2024) or implicitly (Goyal et al., 2024; Wang et al., 2024d; Pfau et al., 2024; Chen et al., 2024). In this work, we require LLMs to explicitly speak with their intent in an iterative and dynamic manner, as shown in Figure 1(a), instead of generating all plans first and then solving the task (Wang et al., 2023; Li & Zhang, 2024). The existing iterative planning framework (Lyu et al., 2024) focuses on addressing information retrieval issues in knowledge-intensive long-form generation tasks, while our SWI method generates intents that function as strategic planning to guide problem analysis and logical reasoning.
Intent-related Research.
Intent Detection (ID) and New Intent Discovery (NID) (Kumar et al., 2022; Liang et al., 2024; Zhang et al., 2024a; Tang et al., 2024; Zhang et al., 2024c; Qian et al., 2024; Yin et al., 2025), which classify utterances into known or novel intent categories, are longstanding challenges in natural language understanding (Larson et al., 2019; Casanueva et al., 2020; Zhang et al., 2021; Weld et al., 2022). Typically, these tasks are approached as classification problems (Wang et al., 2024b; Yoon et al., 2024; Zhang et al., 2024b; Sakurai & Miyao, 2024), where models assign sentences to predefined intent classes. In contrast, Speaking with Intent (SWI) generates intent as free-form text rather than fixed categories, enhancing flexibility and fluency. SWI naturally integrates intent statements as planning into the reasoning process, providing contextual guidance for subsequent analysis.
@@ 3 Speaking with Intent
This section presents the problem-solving workflow of LLMs and introduces Speaking with Intent (SWI), enabling LLMs to explicitly articulate their intent for planning and reasoning.
@@ 3.1 Problem-solving Workflow using LLMs
Let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be a dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information (questions), $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding references (correct answers), and $n$ is the number of instances in $\mathcal{D}$. For mathematical reasoning datasets, $X_{i}$ is the math problem, and $y_{i}$ is the correct answer (usually an integer number). For multiple-choice QA datasets, $X_{i}$ contains the question and options, and $y_{i}$ is the answer label such as (A)/(B)/(C) or Yes/No. For text summarization datasets, $X_{i}$ is the source article, and $y_{i}$ is the reference summary.
@@
3.2 Baseline vs. Speaking with Intent
In this work, we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts $P_{s}$ and restating the SWI requirement in the user prompt $P_{u}$. Figure 2 presents the system prompt and user prompt of Baseline generation ($P_{s}^{\text{base}}$ and $P_{u}^{\text{base}}$) and our SWI method ($P_{s}^{\text{swi}}$ and $P_{u}^{\text{swi}}$).
Figure 2: LLM Prompts. The system and user prompts of Baseline and our SWI method.
@@ 3.3 Result Evaluation
Lastly, we extract the final answer (after “Final Answer:” in $\hat{y}_{i}$), denoted as $\tilde{y}_{i}$, and compute the overall performance of $\mathcal{M}$ on the dataset $\mathcal{D}$ by
@@
For mathematical reasoning tasks, we first extract numbers in $\tilde{y}_{i}$ and apply text normalization to both $\tilde{y}_{i}$ and the reference $y_{i}$, and then conduct exact match to check if the generated answer $\tilde{y}_{i}$ is correct.
For multiple-choice QA tasks, we adopt the Option Selection metric introduced by Yin & Carenini (2025), which evaluates the LLM perplexity of different option concatenations and selects the one with lowest perplexity as the model’s choice.
To evaluate the quality of summaries, we apply the standard ROUGE (Lin, 2004) as the automatic evaluation metric $\mathcal{S}$ and complement it with a more sophisticated fact-checking analysis as described in Section 5.2.
4 Main Experiments on Mathematical Reasoning
This section presents the main experiments to verify the effectiveness of SWI on mathematical reasoning. To ensure reproducibility, we fixed all random seeds, set the generation temperature to zero, and conducted all experiments twice, yielding reproducible results. [^2] Please refer to Appendix A for the reproducibility statement and Appendix B for dataset details.
4.1 Models
@@
In all experiments, we employ LLaMA3-8B-Instruct (Dubey et al., 2024) as the model $\mathcal{M}$ for generation and evaluation. It is an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters. We use the model checkpoint and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020).
4.2 Datasets
The proposed SWI method aims to enhance LLM reasoning, so we conduct extensive experiments on various mathematical reasoning benchmarks, ranging from grade-school to competition-level problems. In the math reasoning task, the model is asked to solve the given math problem and present the final answer. We consider grade-school math problems including GSM8K (Cobbe et al., 2021), GSM8K-Platinum (Vendrow et al., 2025), and MATH500 (Lightman et al., 2024), as well as three competition-level math benchmarks: AMC23, AIME24, and AIME25, which have 40, 30, and 30 math problems, respectively. Despite the smaller scale, these math problems are much more challenging than those in GSM8K and MATH500. The dataset statistics are presented in Table 1.
[…]
Mathematical Reasoning GSM8K (Cobbe et al., 2021) Test 1,319 124
GSM8K-P (Vendrow et al., 2025) Test 1,209 124
MATH500 (Lightman et al., 2024) Test 500 134
[…]
4.3 Experimental Comparison
The main comparison is LLM generation with intent (SWI) or without intent (Baseline), and the effectiveness of SWI is verified if the former outperforms the latter. For mathematical reasoning and multiple-choice QA tasks, we also compare SWI with answer-trigger prompting methods CoT (Kojima et al., 2022) and ARR (Yin & Carenini, 2025), as well as previous LLM planning method Plan-and-Solve (PS) prompting (Wang et al., 2023).

### v3

@@ Abstract
Intent, typically clearly formulated and planned, functions as a cognitive framework for communication and problem-solving. This paper introduces the concept of Speaking with Intent (SWI) in large language models (LLMs), where the explicitly generated intent encapsulates the model’s underlying intention and provides high-level planning to guide subsequent analysis and action. By emulating deliberate and purposeful thoughts in the human mind, SWI is hypothesized to enhance the reasoning capabilities and generation quality of LLMs. Extensive experiments on text summarization, multi-task question answering, and mathematical reasoning benchmarks consistently demonstrate the effectiveness and generalizability of Speaking with Intent over direct generation without explicit intent. Further analysis corroborates the generalizability of SWI under different experimental settings. Moreover, human evaluations verify the coherence, effectiveness, and interpretability of the intent produced by SWI. The promising results in enhancing LLMs with explicit intents pave a new avenue for boosting LLMs’ generation and reasoning abilities with cognitive notions. [^1] Source code: https://github.com/YuweiYin/SWI
SWI: Speaking with Intent in Large Language Models
Yuwei Yin1, EunJeong Hwang1,2, Giuseppe Carenini1 1University of British Columbia; 2Vector Institute for AI {yuweiyin,ejhwang,carenini}@cs.ubc.ca
@@ 1 Introduction
Intent, the goal-oriented intention in our mind (Adams, 1986; Mele, 1989; Mele and Moser, 1994), serves as a critical component in communication and a guiding framework for problem-solving. As illustrated in Figure 1(a), human thinking (Kahneman, 2011) typically follows a structured loop where intent—a mental state or proactive commitment to perform a specific action or produce a particular outcome—directs problem analysis and logical reasoning, therefore also facilitating communication and interaction. Hence, we hypothesize that enabling AI systems to speak with their intent explicitly can replicate this meta-cognitive process, thereby improving their generation quality and reasoning ability.
@@
Figure 1: SWI Overview. (a) The intent, functioning as meta-thought and planning, guides the analysis with reasoning to answer the question. (b) The performance improvements brought by SWI on various tasks.
In recent years, large language models (LLMs) (Zhao et al., 2023; Min et al., 2023; Minaee et al., 2024) have revolutionized Natural Language Processing (NLP) with their excellent generative capabilities (OpenAI, 2024; Anthropic, 2024; Gemini, 2024). Enhancing LLMs on various language understanding and logical reasoning tasks (Hendrycks et al., 2021; Suzgun et al., 2023; Lightman et al., 2024; Vendrow et al., 2025) is vital for their ongoing development (Huang and Chang, 2023; Qiao et al., 2023; Patil, 2025).
@@
This work introduces Speaking with Intent (SWI), requiring LLMs to articulate their own intent as a planning mechanism during generation, which makes LLM intentionality more explicit, in a way reminiscent of the long tradition of intent-driven generation in classic NLG (Grosz and Sidner, 1986; Mann and Thompson, 1988; Moore and Paris, 1993; Reiter and Dale, 2000). In essence, we hypothesize that, due to the autoregressive nature of LLMs (Radford et al., 2019) and the attention mechanism (Vaswani et al., 2017), explicitly stated intent provides high-level guidance for subsequent analysis and reasoning. For example, when applying SWI to summarization tasks, each analytical step in summarizing an article is effectively guided by a preceding intent statement, which is a piece of free-form text generated by instruction-following LLMs (Ouyang et al., 2022; Rafailov et al., 2023; Llama, 2024) instead of a predefined class as in traditional intent modeling (Weld et al., 2022).
@@
In this work, we verify the proposed hypothesis by comprehensively evaluating the effectiveness and generalizability of the proposed SWI method. Specifically, the experimental results across three diverse task categories (i.e., text summarization, multi-task question answering, and mathematical reasoning) demonstrate that speaking with intent in LLMs consistently outperforms directly generating responses without explicit intent. In summarization tasks, SWI produces summaries that are more accurate, concise, and factually reliable, with fewer hallucinations (Ji et al., 2023; Li et al., 2024a) in the output. In math reasoning tasks, SWI surpasses the LLM reasoning method Chain-of-Thought (CoT) (Kojima et al., 2022) and LLM planning method Plan-and-Solve (PS) (Wang et al., 2023), and SWI can work synergistically with these methods to further improve them. Additionally, we perform human evaluations to assess the coherence, effectiveness, and interpretability of the intent generated by our SWI method. Evaluators largely agree on the quality of the generated intent, particularly for mathematical reasoning tasks. The evaluation results confirm that SWI enhances task performance and output explainability.
The key contributions of this work are as follows: ➊ We introduce Speaking with Intent in LLMs, where the generated intent effectively guides problem analysis, logical reasoning, and text generation, boosting performance across various benchmarks. ➋ Extensive experiments and analyses across diverse task types and multiple datasets, including text summarization, multi-task QA, and mathematical reasoning, demonstrate the consistent effectiveness and generalizability of SWI. ➌ Human evaluations validate the coherence, effectiveness, and interpretability of the intent generated by SWI, with our evaluation practice providing standards for assessing freely generated intents.
@@ 2 Speaking with Intent
This section presents the problem-solving workflow of LLMs and introduces Speaking with Intent (SWI), enabling LLMs to explicitly articulate their intent during response generation.
@@ 2.1 Problem-solving Workflow using LLMs
Let $\mathcal{D}=\{\mathcal{X},\mathcal{Y}\}$ be a dataset, where $\mathcal{X}=\{X_{1},X_{2},\dots,X_{n}\}$ is the input information (questions), $\mathcal{Y}=\{y_{1},y_{2},\dots,y_{n}\}$ is the corresponding references (correct answers), and $n$ is the number of instances in $\mathcal{D}$. For text summarization datasets, $X_{i}$ is the source article, and $y_{i}$ is one of the reference summaries. For multiple-choice QA datasets, $X_{i}$ contains the question and options, and $y_{i}$ is the answer label such as (A)/(B)/(C) or Yes/No. For mathematical reasoning datasets, $X_{i}$ is the math problem, and $y_{i}$ is the correct answer (usually an integer number).
@@
2.2 LLM Speaking with Intent
SWI is a novel method that brings a significant cognitive concept (i.e., intent) into LLM generation. SWI is implemented in a simple and reproducible approach, i.e., we require LLMs to speak with intent (SWI) by presenting detailed instructions in the system prompts $P_{s}$ and restating the SWI requirement in the user prompt $P_{u}$. Table 1 presents the SWI instructions for QA and math tasks. For summarization tasks, the fourth requirement in the system prompt is At last, clearly and concisely give your final summary starting with "Final Summary:" and the user prompt becomes Speak with intent and summarize the following article. {{article}}.
As showcased in Figure 2 (math) and Figure 3 (summarization), LLMs that speak with intent articulate their intents during the thinking, reasoning and communication process and then provide the final answer based on the analysis.
System Prompt: You are a helpful assistant who speaks with intent. During generation, follow all the requirements below: 1. Always explicitly state your own intent before speaking each sentence. 2. Each intent statement should explain the sentence that follows. 3. Your intent must start with the "<INTENT>" tag and end with the "</INTENT>" tag. The content within the intent tags must begin with "To" followed by a verb, such as "To accomplish a task." 4. At last, clearly and concisely give your final answer starting with "Final Answer:" User Prompt: Speak with intent and answer the following question. {{question}}
Table 1: SWI instructions for QA and math tasks.
Figure 2: An example of LLMs speaking with intent for mathematical reasoning tasks.
@@ 2.3 Result Evaluation
We extract the final answer (denoted as $\tilde{y}_{i}$) from the model outputs $\hat{y}_{i}$ and compute the overall performance of $\mathcal{M}$ on the dataset $\mathcal{D}$ by
@@
To evaluate the quality of summaries, we apply the standard ROUGE (Lin, 2004) as the automatic evaluation metric $\mathcal{S}$ and complement it with a more sophisticated fact-checking analysis as described in Section 4.1. For multiple-choice QA tasks, we adopt the Option Selection metric introduced by Yin and Carenini (2025), which evaluates the LLM perplexity of different option concatenations and selects the one with lowest perplexity as the model’s choice. For mathematical reasoning tasks, we first extract numbers in $\tilde{y}_{i}$ and apply text normalization to both $\tilde{y}_{i}$ and the reference $y_{i}$, and then conduct exact match to check if the generated answer $\tilde{y}_{i}$ is correct.
3 Experimental Setup
This section presents the experimental setup, including Tasks and Datasets (§ 3.1), Model Settings (§ 3.2), and Baseline Settings (§ 3.3).
3.1 Tasks and Datasets
The proposed SWI method aims to enhance the generation quality and reasoning ability of LLMs. To comprehensively study the effectiveness and generalizability of SWI, we conduct extensive experiments on various challenging benchmarks of three diverse task types: text summarization (Sum), multi-task question answering (QA), and mathematical reasoning (Math). The dataset statistics are presented in Table 2.
Text Summarization (Sum).
We hypothesize that Speaking with Intent benefits natural language generation tasks like summarization, where the generated intent can guide the model in summarizing the source article point by point in an orderly fashion. Hence, we test the effect of SWI on the following text summarization datasets covering different genres: CNN/DailyMail (CDM) (Hermann et al., 2015; See et al., 2017), Extreme summarization (XSum) (Narayan et al., 2018), XL-Sum (Hasan et al., 2021), DialogSum (Chen et al., 2021), and WikiLingua (Ladhak et al., 2020).
Task Dataset Split Size
Sum
CDM Test 11,490
XSum Test 11,334
XL-Sum Test 11,535
DialogSum Test 1,500
WikiLingua Test 3,000
QA
BBH Test 5,511
MMLU Test 13,842
MMLU-Pro Test 12,032
Math
GSM8K Test 1,319
GSM8K-P Test 1,209
MATH500 Test 500
Table 2: Dataset Statistics.
Multi-task Question Answering (QA).
Our SWI method is also evaluated on various multi-task question answering datasets, including BIG-Bench Hard (BBH) (Suzgun et al., 2023), MMLU (Hendrycks et al., 2021), and MMLU-Pro (Wang et al., 2024b). They are all reasoning-intensive benchmarks designed as multiple-choice QA tasks, where the model is asked to select the most appropriate one from the given options to answer the question. Here, the hypothesis is that generating its intent explicitly as text improves the system’s question analysis abilities.
Mathematical Reasoning (Math).
Beyond multiple-choice QA tasks, we also consider mathematical reasoning benchmarks, where the model is asked to solve the given math problem and present the final answer (numerical values). We consider representative and high-quality math benchmarks, including Grade School Math 8K (GSM8K) (Cobbe et al., 2021), GSM8K-Platinum (Vendrow et al., 2025), and MATH500 (Lightman et al., 2024). Again, our hypothesis is that speaking with explicit intent improves the model’s reasoning abilities.
3.2 Model Settings
Language Models.
@@ 3.2 Model Settings
By default, we employ LLaMA3-8B-Instruct (Llama, 2024) as the language model $\mathcal{M}$ for generation and evaluation. It is an open-weights, instruction-following, and Transformer-based (Vaswani et al., 2017) LLM with $8$ billion model parameters. We load the model checkpoint and tokenizer provided by Hugging Face Transformers (Wolf et al., 2020). To further assess the generality of SWI, we also evaluate the efficacy of SWI with different LLMs in § 5.3. In the fact-checking evaluation for summaries (§ 5.1), we adopt GPT-4o-mini (OpenAI, 2024) to decompose the generated summary and reference summary into two sets of atomic facts.
Generation Configurations.
Each experiment session was conducted on a single NVIDIA H100 GPU, and all the models were loaded in a half-precision mode (float16). The input sequence is not truncated to avoid losing context information, while we set the maximum number of newly generated tokens to $4096$ during generation.
Reproducibility Statement.
To guarantee reproducibility, we fixed the seeds to 42 for all random modules, set the LLM generation temperature to 0 for deterministic generation without sampling, and ran all experiments three times, obtaining reproducible generation outputs and evaluation scores. The source code is available on GitHub.
3.3 Baseline Settings
The main comparison is LLM generation with intent or without intent, and the effectiveness of SWI is verified if the former outperforms the latter.